OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 2x + 3y = 11 –4x + 2y = 2 A. (1, 3) B. (2, 3) C. (1, 2) D. (1, –2)

OpenStudy (anonymous):

@DSS Or better yet @Champion (DSS) plz help

OpenStudy (anonymous):

This one is less complicated. 2x + 3y = 11 2(2x+3y=11) 4x+6y=22 –4x + 2y = 2 -4x+2y=2 8y=24 y=3 Get x: 2x+3(3)=11 2x+9=11 2x=2 x=1 (1,3)

OpenStudy (anonymous):

thx :D

OpenStudy (anonymous):

You're welcome. Do you understand?

OpenStudy (anonymous):

yes now i do thx, in fact i even inquired openstudy if they could give u a reward of some sort :D

OpenStudy (anonymous):

Oh, thanks. That's not necessary...I just enjoy helping.

OpenStudy (anonymous):

well u do deserve a reward :D

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

np